OpenStudy (anonymous):

find the derviative of 4-ln(x)/2+ln(x) and its domain.

OpenStudy (anonymous):

this is \[f(x)=\frac{4-\ln(x)}{2+\ln(x)}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so far i have \[2+\ln(x)\times0-1 \div(x)-4-\ln(x)\times1\div(x) all divide by (2+(\ln(x))^{2}\]

OpenStudy (anonymous):

use "quotient rule"

OpenStudy (anonymous):

yes i think you have it. denominator is \[(2+\ln(x))^2\]

OpenStudy (anonymous):

i did I just need help on simplifying

OpenStudy (anonymous):

numerator is \[(2+\ln(x))\times \frac{1}{x}-(4-\ln(x))\times \frac{1}{x}\]

OpenStudy (anonymous):

and the domain

OpenStudy (anonymous):

multiply out in the numerator and you get \[\frac{2+\ln(x)-4+\ln(x)}{x}\]

OpenStudy (anonymous):

oh damn i messed up sorry

OpenStudy (anonymous):

it should be numerator of \[(2+\ln(x))\times -\frac{1}{x}-(4+\ln(x))\times \frac{1}{x}\]

OpenStudy (anonymous):

now multiply out and get \[\frac{-2-\ln(x)-4+\ln(x)}{x}\] \[=\frac{-6}{x}\]

OpenStudy (anonymous):

so your "final answer" is \[-\frac{6}{x(2+\ln(x))^2}\]

OpenStudy (anonymous):

okay i got it now could you explain the domain

OpenStudy (anonymous):

answer is right, but i made yet another typo above should be \[2+\ln(x))\times -\frac{1}{x}-(4-\ln(x))\times \frac{1}{x}\]

OpenStudy (anonymous):

as for the domain you are not allowed to divide by 0

OpenStudy (anonymous):

and you are not allowed to take the log of a negative number

OpenStudy (anonymous):

so you have to make sure that \[\ln(x)+2\neq 0\] \[\ln(x)\neq -2\] \[x\neq e^{-2}\]

OpenStudy (anonymous):

therefore domain is \[(0,e^{-2})\cup (e^{-2},\infty)\] which is a long winded way of saying that x must be positive and not equal to e^-2

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

could you also help me with f(x)=[ ln(x) ]^2 / x^2